OpenStudy (anonymous):

Show an equation and a solution. How much pure antifreeze must be added to 12 L of a 40% solution to obtain a 60% solution. (Remember that pure antifreeze is 100% = 1.)

OpenStudy (jackellyn):

Number of liters of pure antifreeze to be added (x): 0.4(12) + x = 0.6(12 + x) 2(12) + 5x = 3(12 + x) 24 + 5x = 36 + 3x 2x = 12 x = 6 Answer: 6 liters of pure antifreeze must be added.

OpenStudy (jackellyn):

Proof new mixture is 60% solution: = 100%([0.4{12 liters} + 6 liters]/[12 liters + 6 liters]) = 100%([4.8 liters + 6 liters]/18 liters) = 100%(10.8 liters/18 liters) = 100%(0.6) = 60%

OpenStudy (anonymous):

at the moment you have 40% of 12 L = \(.4\times 12=4.8\) L of antifreeze add say \(x\) amount of pure antifreeze, then you have \(4.8+x\) liters of a antifreeze and \(12+x\) liters of solution you want this \(12+x\) liters to be 60% antifreeze, so set \(4.8+x=.6(12+x)\) and solve for \(x\) probably easiest to get rid of the decimals by multiplying both sides by 10 to get \[48+10x=6(12+x)\] \[48+10x=72+6x\] \[4x=24\] \[x=6\]

OpenStudy (anonymous):

same answer and almost same method as jackellyn wrote above, except for line number 2 where she went from \[.4\times 12+x=.6(12+x\] to \[2\times 12+5x=3(12+x)\] where the step was multiplication by 5

OpenStudy (anonymous):

thnks!!!